An aluminium cable has a resistance of 27ohms at a temperature of 35
1
Expert's answer
2021-09-08T17:45:06-0400

"Resistance at \theta ^{^{\circ}}C, R_{\theta }=R_{0}(1+\alpha_{0}\theta )θ"

Hence resistance at 0°C,

"R \n0\n\u200b\t\n = \n\n\\frac{R \n\u03b8}{(1+\u03b1 \n0\n\u200b\t\n \u03b8)}\n\u200b\t\n \n\u200b" = "\\frac{27}{[1+(0.0038)(35)]}\n\n\u200b"

"= \frac{27}{1.33}" = 20.3 ohms